# reference.bib

### v1

@@ IV EXPERIMENTS
In this section, we will discuss in detail the setup of the Formula 1 tracks domains, followed by the MPC design in this testing domain. Additionally, we will present experimental results from both training and holdout tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
@@ IV-A Experiment Setup
The experiments are done in high-speed Formula 1 tracks [f1tenth_racetracks] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively. The reference trajectory is represented as a set of waypoints in the center of the track. The track size is downscaled to a 10:1 ratio to make each lap a reasonable length. The reference speed for the vehicle in the expert MPC is $10m/s$. Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track. The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions. The path tracking is challenging for traditional gradient-free MPC because the vehicle is running at a high speed and the tracks have multiple sharp turns. In this scenario, traditional gradient-free MPC solvers could not optimize the control solution in real time, which makes it unreasonable for real-world application.
@@
As shown in Figure 2, without a specified threshold, optimization continues until reaching the maximum iteration limit. In our experiments, we found that the MPC cost reduces subtly during the later optimization iterations. As such, it is important to provide the solver with an early-stop criterion. We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$. Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.08 seconds in worst-case scenarios. Details of the MPC design and hyper parameters in MPC and training are presented in Appendix VI-A and VI-B
(a) Training tracks
(b) Zero-shot tracks (complex)
(c) Zero-shot tracks (simple)
Figure 3: Tracks tested during training and testing.
IV-B Experiment Results
During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step. We benchmark our algorithm’s performance against MPC using all-zero initial guesses and initial guesses derived from the MPC’s solution at the previous step. We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)). The results are shown in Figure 4 and Figure 5, respectively.
IV-B1 Optimization Time
On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time, falling well within the upper bound of $0.08s$. Additionally, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time compared to the policy trained solely through offline behavior cloning, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality problem caused by behavior cloning (BC). Furthermore, our two-phase training framework achieves more significant improvement in the zero-shot domains, indicating that the online fine-tuning phase alleviates the covariance shift problem caused by BC and enhances the model’s generalizability.
IV-B2 Control Performance ($xte$)
In the context of the training tracks, the warm-start policy trained through offline and online fine-tuning outperforms the policy trained solely through offline behavior cloning in terms of tracking accuracy, further illustrating the effectiveness of our online fine-tuning algorithm in addressing the suboptimality problem caused by BC. Given that the width of the downscaled track is $2.2m$, achieving an $xte$ of less than $0.3m$ demonstrates precise tracking of the reference path. Additionally, our two-phase training framework exhibits more significant improvement in the zero-shot domains, reinforcing the efficacy of the online fine-tuning phase in mitigating the covariance shift problem caused by BC and improving the model’s generalizability.
Figure 4: Experiment results on training tracks.
Figure 5: Experiment results on the six zero-shot testing tracks.
The results also reveal that employing an MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane and struggles to navigate sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
As the real-time MPC with either an all-zero initial guess or an initial guess derived from the MPC’s previous solution fails to complete laps on both the training and challenging zero-shot tracks, we conducted further testing on four warm start policies on a simple zero-shot track, IMS, as depicted in Figure 3(c). The results are summarized in Figure 6. Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
Figure 6: Experiment results on the simple IMS track.
As the vehicle route on the track, we further compute the curvature of the track surrounding the vehicle. This analysis allows us to illustrate the correlation between the curvature and lateral error ($xte$), depicted in Figure 7. It shows that our second phase fine tuning has an advantage on average $xte$ and this advantage has a modest increase when the track is less curvy. This outcome underscores the efficacy of our proposed algorithm in enhancing the precision of the planned trajectory. Details regarding the method employed to compute the curvature of the track around the vehicle are provided in the Appendix VI-C
Figure 7: We run the MPC with two types of warm-start policy (BC and ours) on the ”Catalunya” track, where we gather data on the $xte$ and track curvature surrounding the vehicle. Subsequently, we depict these data points in the figure above.
In summary, our empirical results support that:
  • MPC with initial guesses trained by data-driven methods significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
  • The two-phase training shows superior performance in zero-shot scenarios, indicating better generalizability.
V LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are several limitations and avenues for future research that merit consideration. One limitation of our current experiments is that we have only focused on autonomous driving where the action space is limited to two dimensions (i.e., steering and acceleration). Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces. Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI. Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [williams2015model]. MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost. Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution. This reduces the iteration count needed for optimization, aligning with early stopping criteria. While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
VI CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase learning framework combines offline behavior cloning and online fine-tuning to provide improved initial guesses for the MPC solver. Experimental results on both training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and enhancing path tracking accuracy. Our learning framework integrated with MPC opens up new avenues for improving the efficiency and applicability of trajectory optimization in various dynamic systems.
APPENDIX
VI-A MPC Design Details
@@
The MPC objective function is composed of five parts as shown in Equation (7). The first two terms are $xte$ and Error in Heading ($eth$). $xte$ is computed using Equation 4. $eth$ denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. And $eth$ is computed using Equation (8). $v_{t}^{\text{ref}}$ is the desired speed of the vehicle. The last two terms in Equation (7) regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead.
@@
VI-B Hyper Parameters in MPC and Training
Hyperparameters Values
Weight in MPC cost function $w_{0}$ 2000
Weight in MPC cost function $w_{1}$ 100
Weight in MPC cost function $w_{2}$ 60
Weight in MPC cost function $w_{3}$ 2
Weight in MPC cost function $w_{4}$ 20
Weight in MPC cost function $w_{5}$ 300
Loss weight coefficient $\lambda$ 0.9
Loss weight coefficient $\lambda_{1}$ 0.5
Loss weight coefficient $\lambda_{2}$ 0
Loss weight coefficient $\lambda_{3}$ 0.5
MPC planning horizon 25
Length of each time step 0.02 second
Reference speed $10m/s$
Wheelbase 2.89 m
Expert MPC maximum optimization iteration 300
Real-time MPC maximum optimization iteration 300
Real time MPC early stop criteria $xte<0.1$
VI-C Curvature Computation
Algorithm 3 Curvature Computation
1:Extract 10 consecutive nearest waypoints ahead of the vehicle, $\mathbf{P}=\{\mathbf{p}_{1},\mathbf{p}_{2},...,\mathbf{p}_{10}\}$. $\mathbf{p}_{i}=(x_{i},y_{i})$ represent the $i$-th waypoint
2:for $\mathbf{p}_{i-1}$, $\mathbf{p}_{i}$, and $\mathbf{p}_{i+1}$ in $\mathbf{P}$ do
3:     Compute the vector $\mathbf{v}_{1}=\mathbf{p}_{i}-\mathbf{p}_{i-1}$
4:     Compute the vector $\mathbf{v}_{2}=\mathbf{p}_{i+1}-\mathbf{p}_{i}$
5:     Calculate the dot product $\mathbf{v}_{1}\cdot\mathbf{v}_{2}$
6:     Calculate the magnitudes $\|\mathbf{v}_{1}\|$ and $\|\mathbf{v}_{2}\|$
7:     Compute the $\cos$ of the angle between the vectors:
$\cos(\theta_{i})=\left(\frac{\mathbf{v}_{1}\cdot\mathbf{v}_{2}}{\|\mathbf{v}_{1}\|\|\mathbf{v}_{2}\|}\right)$
8:end for
9:$Curvature=\sum_{i=2}^{9}\left(1-cos(\theta_{i})\right)$

### v2

@@ IV EXPERIMENTS
In this section, we will discuss in detail the setup of the Formula 1 track domains, followed by the MPC design in this testing domain. Additionally, we will present experimental results from both training and zero-shot tracks, demonstrating how our algorithm enhances MPC’s performance in terms of both optimization time and tracking accuracy.
@@ IV-A Experiment Setup
The experiments are done in high-speed Formula 1 tracks [26] which are divided into three training tracks and seven zero-shot testing tracks, as shown in Figure 3(a), Figure 3(b), and Figure 3(c), respectively. The reference trajectory is represented as a set of waypoints in the center of the track. The track size is downscaled to a 10:1 ratio to make each lap a reasonable length. The reference speed for the vehicle in the expert MPC is $10m/s$. Given the downscaled size of the track data, the vehicle could actually speed up to $100m/s$ in the 1:1 scale track. The friction between the tire and the road is not considered in our dynamics model, as it depends on multiple factors such as tire pressure, temperature, humidity, and road conditions.
@@
The MPC objective function is composed of five parts as shown in Equation (7). The first two terms are $xte$ and Error in Heading ($eth$). $xte$ is computed using Equation 5. $eth$ is computed using Equation (8), which denotes the angular disparity between the intended path direction and the current heading of a vehicle in path tracking systems. $v_{t}^{\text{ref}}$ is the desired speed of the vehicle. The last two terms regulate the rate of change of the steering angle and acceleration to make planned trajectory smoother. $w_{0},w_{1},w_{2},w_{3},w_{4}$ are the coefficients balancing the importance of each term. The planning horizon of the MPC is 25 steps and the planning step $dt$ is 0.02 seconds, which means that the MPC looks 0.5 seconds ahead. Since the calculation of $xte$ involves finding the closest point on a reference trajectory to the car’s current position, it introduces discontinuities and non-differentiability into the MPC objective function. As a result, only gradient-free, sampling-based solvers are suitable for this problem [27, 28]. Here we choose to use the gradient-free solver COBYLA [29] in our experiments.
Furthermore, given the high speed of the vehicle and the presence of multiple sharp turns on the track, gradient-free MPC solvers without a good initial guess of the control solution will struggle to optimize the trajectory in real time, making real-world implementation challenging. And we show in Section IV-C gradient-free MPC solver only succeeds in simple IMS track and fails at all of the complex tracks.
@@
We implement the early-stop condition based on a planned trajectory’s accumulated cross-track error ($xte$) being $<0.1m$ to avoid MPC wasting time on optimizing a trajectory that is already good enough. Additionally, to ensure real-time planning during testing, we cap the maximum optimization iterations at 50, which guarantees that the optimization time is less than 0.05 seconds in worst-case scenarios. As, 20Hz running rate of MPC is validated by [30] for vehicle path tracking.
(a) Training tracks
(b) Zero-shot tracks (complex)
(c) Zero-shot tracks (simple)
Figure 3: Tracks tested during training and testing.
IV-B Baselines and Metrics
During testing, we evaluate on two metrics: the average MPC optimization time (seconds) per step and the average $xte$ (m) per step. We compare our algorithm’s performance against five types of initial guess policies: All Zero: all-zero initial guesses, Previous Solution: initial guesses derived from the MPC solution at the previous step, Ours w/o BC: initial guesses from a warm-start policy trained exclusively with an online training algorithm, without offline training, Ours w/o RL: initial guesses from a warm-start policy trained solely with offline BC, and Ours w/o DAgger: initial guesses from a warm-start policy trained with our two-phase learning algorithm but without the use of DAgger. To ensure a fair comparison between the Ours w/o BC and warm-start policy trained by our two-phase learning algorithm, we extend the training time of Ours w/o BC by an additional hour to account for the data collection time in the offline BC phase.
IV-C Experiment Results
We perform testing on both the training tracks (Figure 3(a)) and the challenging zero-shot tracks (Figure 3(b)). The results are shown in Figure 4 and Figure 5, respectively. On both the training and zero-shot tracks, employing a warm-start policy trained through either offline learning or a combination of offline and online fine-tuning significantly reduces MPC optimization time and improves tracking accuracy. Besides, the warm-start policy trained with both offline and online fine-tuning achieves better MPC optimization time and tracking accuracy compared to the policy trained solely through offline BC, demonstrating the capability of our online fine-tuning algorithm in addressing the suboptimality covariance shift problem caused by BC.
@@ IV-C Experiment Results
The results also reveal that initialize an MPC with either All Zeros or Previous Solution fails to complete laps on both the training and zero-shot tracks. The vehicle consistently deviates from the lane when facing any sharp turns. This limitation arises because the MPC often requires more iterations to optimize the control solution as the vehicle approaches the curves of the track. Since the real-time MPC only optimizes the solution for a maximum of 50 iterations at each step, the returned solution lacks the optimization necessary to guide the vehicle through sharp turns effectively. This underscores the necessity of a well-informed initial guess to minimize the number of optimization iterations required.
Figure 4: Experiment results on training tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Figure 5: Experiment results on the six zero-shot testing tracks. The asterisk (*) implies statistical significance: * $p<0.05$, ** $p<0.01$, *** $p<0.001$, n.s. means no significant.
Furthermore, the Ours w/o RL algorithm also fails on both training and zero-shot domains. This is because the absence of offline BC requires the algorithm to explore the solution space from scratch, causing the online learning algorithm to take significantly longer to converge to a satisfactory solution. Given the same amount of training time, the quality of the policy learned by Ours w/o RL lags behind that of the policy learned through the two-phase training algorithm. This demonstrates the effectiveness of offline BC in accelerating online training.
To further inspect the relative performance of our approach when our baselines performed better, i.e. on simpler tracks, we conducted further testing on six warm-start policies on a simpler zero-shot track, IMS, as depicted in Figure 3(c),   similar to the path-tracking scenario in [31]. The results are summarized in Figure 6. Our method demonstrates a significant reduction in MPC optimization time while maintaining high tracking accuracy, showcasing its effectiveness in enhancing MPC performance.
Figure 6: Experiment results on the simple IMS track.
In summary, our empirical results support that:
  • MPC with warm-start policy trained by either offline learning or a combination of offline and online fine-tuning significantly outperforms those with all-zero or previous solution-derived guesses, crucial for navigating sharp turns.
  • The offline training algorithm significantly warm-starts the online training algorithm and make it take much less training time to get a satisfying performance.
  • The two-phase training algorithm outperforms in zero-shot scenarios, demonstrating superior generalizability.
V LIMITATIONS AND FUTURE WORKS
While our proposed two-phase learning framework shows promising results in expediting optimization processes and enhancing control performance for robot control tasks, there are still limitations and avenues for future research that merit consideration. Future work could expand into more complex domains, such as robot arm manipulation or quadcopter control, to explore the algorithm’s efficacy in higher-dimensional spaces. Additionally, future research could also involve utilizing our proposed algorithm to expedite other iterative control optimization techniques like MPPI. Although our algorithm primarily interfaces with MPC, it is versatile enough to serve as a warm start for other iterative control optimization methods like Model Predictive Path Integral (MPPI) control [32]. MPPI typically employs stochastic optimization, initializing with a guess of the control sequence distribution, sampling sequences, and iteratively refining the distribution to minimize control cost. Our framework can be trained to predict the trajectory Gaussian distribution mean, enhancing MPPI’s performance by offering a superior initial distribution. This reduces the iteration count needed for optimization, aligning with early stopping criteria. While MPC suits systems with complex dynamics and constraints, we choose it as our focus due to its representativeness in trajectory optimization methods.
VI CONCLUSIONS
In this paper, we introduce a novel approach to accelerate MPC optimization by learning a warm-start policy. Our two-phase framework combines offline BC and online fine-tuning to provide better initial guesses for the MPC solver. Experiments on training and zero-shot tracks demonstrate the effectiveness of our approach in reducing optimization time and improving path tracking accuracy. This integration of learning with MPC enhances the efficiency and applicability of trajectory optimization in dynamic systems.